OpenStudy (anonymous):

Rewrite in sigma notation: 5+8+11+14+...+50.

OpenStudy (anonymous):

50 = 5 + (n-1) 3 45 = 3n - 3 3n = 48 n =16

OpenStudy (anonymous):

5 + 3n -3 = 2 + 3n

OpenStudy (anonymous):

i hope this helps

OpenStudy (lgbasallote):

i thought sigma notation had..i dont know..a sigma? \(\Sigma\)

OpenStudy (anonymous):